9x^2-120x-81=0
How is this worked out?

Answer by checkley75(3666)

9X^2-120X-81=0
USING THE QUADRATIC EQUATION WE GET:
X=(120+-SQRT[-120^2-4*9*-81])/(2*9)
X=(120+-SQRT[14400+2906])/18
X=(120+-SQRT17316)/18
X=(120+-131.59)/18
X=(120+131.59)/18
X=251.59/18
X=13.977 ANSWER
X=(120-131.59)/18
X=-11.59/18
X=-.644 ANSWER.